OpenStudy (anonymous):

Solve for x. 2(–2 + 2x) = 4(x – 3)

OpenStudy (anonymous):

ok since in the ()'s there not the same you would have to take 2 of 29-2+2x) and multiply it like 2 x -2 =

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

2(-2+2x) *

OpenStudy (anonymous):

and hello :-)

OpenStudy (anonymous):

There is no solution. -4+4 x=-12+4 x

OpenStudy (anonymous):

so -4+4x?

OpenStudy (anonymous):

Subtract 4x from each side. One is left with -4 = -12 which is not true.

OpenStudy (anonymous):

Oh i think I understand

OpenStudy (anonymous):

yea basically

OpenStudy (anonymous):

Yes! thank you to both of you. Since I can only give one emdal I will give medal to shorty and fan rob

OpenStudy (anonymous):

yay thanks @thebossman